OpenStudy (anonymous):

tell wether the equation is a solution to the linear equation (0,2);3x-2y=4,-2x-y=-2

OpenStudy (anonymous):

To tell whether this solution works, you must plug (0, 2) into both equations. Put a 0 where you see an x and a 2 where you see a y. 3(0) - 2(2) = -4 -2(0) - 2 = -2 Since -4 does not equal 4 in the first equation, (0, 2) is not a correct solution, even though (0, 2) did work for the other equation.